OpenStudy (anonymous):

permutations : teacher must arrange 5 tunes for band. she has 20 tunes to choose from. how many arrangements are possible? hint: it's not 20 factorial

OpenStudy (anonymous):

didn't we just do this one?

OpenStudy (anonymous):

it is not 20 factorial because there are only 5 tunes

OpenStudy (anonymous):

did you compute \[20\times 19\times 18\times 17\times 16\]?

OpenStudy (anonymous):

It is a combination, presuming that she does not use the same song more than once. 20!/5!15! is the answer.

OpenStudy (anonymous):

@EulersEquation it says "arrangements" so i assume order matters

OpenStudy (anonymous):

"arrangments" means "songs."

OpenStudy (anonymous):

that, and the fact that the heading says "permutations"

OpenStudy (anonymous):

i got it 20!/ (20-5)!

OpenStudy (anonymous):

She has 20 to choose from. Assuming that she chooses 5 different songs then order does matter. In other words, ABCDE is the same as EDCBA.

OpenStudy (anonymous):

no i don't think so i think "arrangements" means the number of ways to arrange the songs so "put a ring on it" then "louie louie" counts different then the other way around

OpenStudy (anonymous):

@soapia what you wrote is identical to what i wrote

OpenStudy (anonymous):

Well, that needs to be clarified, because as a musician, the term "arrangement" simply refers to "a song."

OpenStudy (anonymous):

not unless you are an arranger

OpenStudy (anonymous):

I would presume that the problem means the same thing. AB and BA are the same arrangement of songs. If it is the answer is 20!/5!/15!. Otherwise, the answer is 20!/5!.

OpenStudy (anonymous):

I am a 17th century mathematician, and I invented most of the math that is in use today; yet I don't get no respect!